[figure: 9.7 cm]

ELECTRONIC, MAGNETIC AND TRANSPORT PROPERTIES

OF YCo_{3}B_{2} COMPOUND

A. Kowalczyk, A. Jezierski, L. Smardz, J. Baszyński

Institute of Molecular Physics, Polish Academy of Sciences, Smoluchowskiego 17, 60-179 Poznań, Poland

We have studied the electronic, magnetic and transport properties of the hexagonal YCo_{3}B_{2} compound. The electronic structure has been studied by X-ray photoemission spectroscopy (XPS) and ab initio self-consistent tight binding linear muffin tin orbital (TB LMTO) method. The XPS spectra are compared with the calculated electronic density of states. Theoretical calculations showed that the YCo_{3}B_{2} is a paramagnet in agreement with experimental results. The main magnetic contribution to the density of states is from Co atoms. Electrical resistivity at low temperatures shows a T^{2} dependence, implying that the scattering by spin fluctuations is dominant in this temperature range. The results obtained for YCo_{3}B_{2} are compared with properties of YCo_{5} compound.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : A. Kowalczyk

Address for correspondence : Institute of Molecular Physics, Polish Academy of Sciences, Smoluchowskiego 17, 60-179 Poznań, POLAND

Email address : ankow@ifmpan.poznan.pl